1. When ONLY Total Costs (TC) are known, explain how to calculate each of the following:

a. Fixed Costs (FC)

b. Variable Costs (VC)

c. Average Variable Costs (AVC)

d. Average Total Costs (ATC)

e. Average Fixed Costs (AFC)

f. Marginal Costs (MC)
Answer #1

a)

Fixed Cost is equal to TC at zero output. Look for TC at zero output,

FC=TC at output=0

It remains the same for all output levels.

b)

Variable Cost is equal to the difference between TC and VC. So,

Variable Cost (VC)=TC-FC

c)

Average Variable Cost is equal to variable cost divided by output level.

AVC=VC/Q (Q=output)

d)

Average Total Cost is equal to total cost divided by output level.

ATC=TC/Q (Q=output)

e)

Average Fixed Cost is equal to fixed cost divided by output level.

AFC=FC/Q (Q=output)

f)

Marginal Cost=Change in TC/Change in Q